# DSwinIR: Rethinking Window-based Attention for Image Restoration

### v2

@@ IV-B2 Robustness in Deweathering Scenarios
On the comprehensive AllWeather dataset collection, which includes snow, raindrops, and outdoor rain, DSwinIR demonstrates clear superiority. As presented in Tab. IV, our model achieves an average PSNR of 34.08 dB, surpassing the previous best, Histoformer [104], by a convincing margin of 0.4 dB. The visual results in Fig. 5 are even more compelling. Compared to prior works, including diffusion-based models, DSwinIR consistently produces images with better clarity, more vibrant colors, and sharper details, effectively removing weather artifacts without introducing secondary distortions.
TABLE V: Comparison on de-weathering tasks on real-world datasets following [85]
@@
SPA+ Chen et al. [98] 37.32 0.97
WGWSNet [85] 38.94 0.98
TransWeather [94] 33.64 0.93
@@
RealSnow Chen et al. [98] 29.37 0.88
WGWSNet [85] 29.46 0.85
TransWeather [94] 29.16 0.82
DSwinIR (Ours) 33.80 0.93
REVIDE Chen et al. [98] 20.10 0.85
WGWSNet [85] 20.44 0.87
TransWeather [94] 17.33 0.82
@@
Average Chen et al. [98] 28.93 0.90
WGWSNet [85] 29.61 0.90
TransWeather [94] 26.71 0.86
DSwinIR (Ours) 34.85 0.93
@@
The true test of a deweathering model lies in its performance on real-world data. On a challenging benchmark composed of real-world rain, snow, and haze datasets [85], DSwinIR again proves its mettle. As shown in Tab. V, our model achieves an average PSNR of 34.85 dB, significantly outperforming the previous state-of-the-art WGWSNet [85] by 0.84 dB. This robust performance on both synthetic and real-world data underscores the practical applicability and strong generalization of our proposed architecture for tackling adverse weather conditions.
TABLE VI: Quantitative results on the Composite Degradation Setting. Our results are highlighted in bold, and the best results are underlined.
@@
Restormer [45] CVPR’22 26.29 28.35 33.10 33.43 24.80 25.28 24.99 26.80 26.15 23.90 23.82 26.99
NAFNet [34] ECCV’22 24.50 25.34 29.24 29.54 21.91 22.75 22.79 23.67 23.86 21.03 20.82 24.13
AirNet[61] CVPR’22 24.83 24.21 26.55 26.79 23.23 22.82 23.29 22.21 23.29 21.80 22.24 23.75
PromptIR[15] NeurIPS’23 26.32 26.10 31.56 31.53 24.49 25.05 24.51 24.54 23.70 23.74 23.33 25.90
TransWeather[94] CVPR’22 23.39 23.95 26.69 25.74 22.24 22.62 21.80 23.10 22.34 21.55 21.01 23.13
WeatherDiff[85] TPAMI’23 23.58 21.99 24.85 24.80 21.83 22.69 22.12 21.25 21.99 21.23 21.04 22.49
WGWSNet[85] CVPR’23 24.39 27.90 33.15 34.43 24.27 25.06 24.60 27.23 27.65 23.90 23.97 26.96
InstructIR[16] ECCV’24 26.70 32.61 33.51 34.45 24.36 25.41 25.63 28.80 29.64 24.84 24.32 28.21
OneRestore[95] ECCV’24 26.55 32.71 33.48 34.50 26.15 25.83 25.56 30.27 30.46 25.18 25.28 28.47
@@
IV-B3 Mastery in Complex Composite Degradations
The ultimate test of a restoration model’s versatility and power lies in its ability to address composite degradations, where multiple types of corruption are simultaneously present. We evaluate DSwinIR on the challenging composite degradation benchmark, which features 11 combinations of low-light, haze, rain, and snow based on CDD-11 dataset [95]. The quantitative results, presented in Tab. VI, unequivocally demonstrate the dominance of our approach. DSwinIR achieves a new state-of-the-art average PSNR of 29.43 dB, significantly surpassing the next best method, OneRestore [95], by a remarkable margin of 0.98 dB. Even more impressively, this superiority is not merely an average tendency; DSwinIR secures the top performance across all 11 individual and combined degradation settings. This consistent, commanding lead is a powerful testament to the robustness and advanced capabilities of our architecture.
In the face of composite degradations like haze and rain, standard attention mechanisms with fixed receptive fields may struggle, either over-smoothing details to remove haze or failing to capture the fine, non-local structure of rain streaks. In contrast, the deformable component of DSwin attention allows the model to dynamically tailor its receptive fields, effectively creating task-specific local experts that can, for instance, focus on thin streak-like patterns for deraining while simultaneously attending to broader regions to model and remove haze. Concurrently, the sliding window mechanism ensures that this powerful local processing is applied efficiently across the entire feature map, preserving spatial context and preventing the loss of fine-grained texture that can occur when dealing with severe, multi-faceted degradation. This synergy allows DSwinIR to decompose and resolve the complex restoration problem without catastrophic interference between sub-tasks.
These quantitative findings are further substantiated by the qualitative results shown in Fig. 6. The visual examples highlight DSwinIR’s superior ability to restore clarity and detail in images afflicted by multiple interacting degradations. Our model more effectively removes all traces of weather artifacts while simultaneously enhancing brightness and contrast, achieving perceptually more natural and artifact-free results. This confirms that DSwinIR’s architectural advantages translate directly into visually compelling and high-fidelity image restoration.
IV-B4 Specialization Power in Single-Task Settings
A defining characteristic of a truly superior architecture is its ability to excel not only as a generalist but also as a specialist. To investigate this “specialization power”, we adapt DSwinIR for single-task training on standard deraining and dehazing benchmarks. These experiments are critical to demonstrate that the performance of our foundational design is not diluted but can be concentrated to achieve mastery, rivaling and even surpassing models designed exclusively for these tasks.
@@ IV-B4 Specialization Power in Single-Task Settings
We begin with the standard Rain100L benchmark. As shown in Tab. VII, the field is highly competitive. When trained solely for this task, DSwinIR establishes a new state-of-the-art result. The performance leap over strong backbone architectures like Restormer (+1.98 dB) is substantial, reaffirming the architectural progress our DSwin attention provides. More tellingly, DSwinIR also narrowly surpasses the latest top-performing model, AdaIR, by 0.12 dB. In a mature benchmark where gains are often incremental, achieving the top position provides unequivocal evidence of DSwinIR’s high-fidelity reconstruction capabilities when fully specialized.
TABLE VII: Single-task deraining results on the Rain100L dataset. DSwinIR establishes a new state-of-the-art, demonstrating its high specialization potential.
Method DIDMDN [105] UMR [106] SIRR [107] MSPFN[26] LPNet[108] AirNet[61] Restormer[45] PromptIR[15] AdaIR[63] DSwinIR (Ours)
@@
The task of single-image dehazing presents a different challenge, often requiring a nuanced understanding of global atmospheric light and local object details. Here, the advantages of DSwinIR’s design become profoundly apparent. As detailed in Tab. VIII, DSwinIR achieves a PSNR of 32.89 dB. This is not merely an incremental improvement but a massive leap in performance, surpassing the powerful AdaIR by an impressive 1.09 dB. This substantial margin is a direct testament to the efficacy of our DSwin attention. We posit that the deformable mechanism’s ability to adapt its receptive field to non-homogenous haze, combined with the sliding mechanism’s capacity to preserve fine structural details, creates a perfect synergy for this task. This result strongly validates that our architectural principles are particularly well-suited for tackling complex, spatially-variant degradations.
TABLE VIII: Single-task dehazing results on the SOTS outdoor test set. DSwinIR demonstrates a significant performance leap over all prior methods.
Method DehazeNet [21] MSCNN [109] AODNet [110] EPDN [111] FDGAN [112] AirNet [61] Restormer [45] PromptIR [15] AdaIR [63] DSwinIR (Ours)
@@
Finally, we test DSwinIR’s specialization power against the ultimate challenge: complex, non-uniform artifacts in real-world images. We use the SPA-Data benchmark, known for its realistic and challenging rain streak patterns. The results, summarized in Tab. IX, are striking. DSwinIR achieves a PSNR of 49.19 dB, establishing a new state-of-the-art. Most notably, it outperforms the highly-optimized and specialized deraining model, DRSformer, by a significant margin of 0.66 dB. Thriving in such a complex, real-world environment, where degradation patterns defy simple synthetic models, is powerful proof of the robustness and superior modeling capacity of the DSwin attention mechanism. It validates that our model’s ability to learn adaptive, data-driven receptive fields is not just a theoretical advantage but a practical necessity for tackling real-world image restoration.
TABLE IX: Single-task deraining results on the real-world SPA-Data dataset. DSwinIR significantly outperforms even highly specialized deraining models.
Method MSPFN [26] MPRNet [33] DualGCN [113] SPDNet [114] Uformer [10] Restormer [45] IDT [40] DRSformer [13] DSwinIR (Ours)
@@
TABLE X: Efficiency evaluation of different methods with 256$\times$256 input.
@@
AirNet [61] Deformable Conv 9M 301G 0.57s 31.20
Uformer [10] Window Attention 20M 41G 0.45s 30.66
Restormer [45] Transposed Self-Attention 26M 155G 0.67s 30.75
PromptIR [15] Trans. SA + Prompts 36M 173G 0.71s 32.06
DRSformer [13] Sparse Window Attention 33M 243G 0.82s 31.26
@@
Figure 7: Ablation studies demonstrating the effectiveness of our key components. We evaluate (1) different attention mechanisms, showing improvements from Sliding Window (31.79 dB) and Deformable Window (32.11 dB) over baselines; (2) DSwin configurations with various kernel sizes ($K$=5, 7, 9) and Multi-scale enhancement (32.69 dB); and (3) FFN module with MSG enhancement, achieving the best performance (32.73 dB). All experiments report the average PSNR of three distinct degradation tasks.
@@ IV-C Ablation Study
To validate our design choices and rigorously quantify the contribution of each proposed component, we conduct a series of in-depth ablation studies. We use the challenging three-task all-in-one setting as our testbed and present the comprehensive results in Fig. 7. Our analysis proceeds methodically, first deconstructing the core DSwin attention mechanism and then justifying the configuration of the full DSwinIR block.
@@
The central thesis of our work is that simultaneously resolving the boundary limitations and the static receptive field issue is key to unlocking superior performance. To prove this, we begin with strong backbone baselines and incrementally introduce our core ideas. The Uformer [10] and DRSformer [13] baselines establish a performance benchmark of 30.66 dB and 31.26 dB, respectively. Our first experiment, isolating the sliding window mechanism, yields a performance of 31.79 dB. This substantial gain over a comparable backbone structure confirms that simply enabling smooth information flow across all spatial locations by adopting a token-centric paradigm provides a powerful performance uplift. In a parallel experiment, isolating the deformable window mechanism boosts the PSNR to an even higher 32.11 dB, even surpassing the prompt-based PromptIR baseline. This result underscores the immense value of a content-adaptive receptive field that can focus on salient features. The true power of our design is revealed when these two concepts are combined. The resulting DSwin module (e.g., with $K$=7) achieves 32.53 dB, a gain that is markedly greater than the sum of the individual improvements. This demonstrates a powerful synergistic effect: the sliding mechanism provides robust structural context, which in turn provides a better feature foundation for the deformable mechanism to learn meaningful semantic offsets.
Figure 8: Visualization of receptive fields of multiple backbones by LAM[115].
@@
The visualizations of the receptive field in Fig. 8 offer profound qualitative insights into why DSwinIR is so effective. Its receptive field is demonstrably both global in reach and local in precision. This is the direct result of our design principles: the sliding mechanism ensures a connected, expansive base, while the deformable mechanism learns to dynamically concentrate attention on areas like object contours and textures.
Moreover, DSwinIR also exhibits exceptional computational efficiency, as detailed in Tab. X. It is crucial to note that DSwinIR achieves its state-of-the-art performance not through brute force, but through superior architectural design. Compared to its top-performing competitors, DSwinIR is significantly more economical. For instance, it requires 15% fewer FLOPs and is 22% faster than PromptIR, while delivering a substantially higher PSNR. The advantage is even more dramatic when compared to DRSformer, against which DSwinIR uses a staggering 40% fewer FLOPs. This potent combination of a more intelligent mechanism and superior computational efficiency establishes DSwinIR not merely as another high-performing model, but as a fundamentally more advanced and practical architecture for the field of image restoration.
@@
Having validated the DSwin concept, we next investigate its optimal configuration. We first analyze the impact of a fixed kernel size, testing $K$=5, $K$=7, and $K$=9. The results show a clear trend: performance improves from 32.02 dB ($K$=5) to 32.53 dB ($K$=7), and then begins to plateau at 32.61 dB ($K$=9). This suggests a trade-off, where a larger receptive field is beneficial, but excessively large kernels may not provide proportionally greater returns for their computational cost. This observation motivates our multi-scale strategy. By employing multi-scale kernels within a single attention module (+Multi-Scale K), we reach a PSNR of 32.69 dB. This result is superior to even the best-performing single-kernel model ($K$=9), providing clear evidence that empowering the model to simultaneously perceive features at multiple scales is more effective than relying on a single, large-but-fixed receptive field.
@@
The final step in our analysis is to justify the design of the feed-forward network. By incorporating our proposed multi-scale gated FFN (+MSG), the model’s performance reaches its peak at 32.73 dB. While the incremental gain at this stage is modest, its importance is strategic. It confirms that maintaining a holistic multi-scale philosophy across the entire Transformer block—in both the attention module and the FFN—is beneficial. This creates a cohesive and synergistic architecture where features are consistently processed and refined across a spectrum of spatial extents, leading to the final state-of-the-art performance. The combination of all these validated components results in our final DSwinIR model.

### v3

@@ IV-B2 Results on Deweathering Datasets
On the comprehensive AllWeather dataset collection, which includes snow, raindrops, and outdoor rain, DSwinIR demonstrates clear superiority. As presented in Table IV, our model achieves an average PSNR of 34.08 dB, surpassing the previous best, Histoformer [Histoformer_SunRGWC24], by a convincing margin of 0.4 dB. The visual results in Figure 6 are consistent with the quantitative results. Compared to prior works, including diffusion-based models, DSwinIR consistently produces images with better clarity, more vibrant colors, and sharper details, effectively removing weather artifacts without introducing secondary distortions.
TABLE V: Comparison on de-weathering tasks on real-world datasets following [weather_data]
@@
SPA+ Chen et al. [ChenHTYDK22] 37.32 0.97
TransWeather [Transweather] 33.64 0.93
WGWSNet [weather_data] 38.94 0.98
@@
RealSnow Chen et al. [ChenHTYDK22] 29.37 0.88
TransWeather [Transweather] 29.16 0.82
WGWSNet [weather_data] 29.46 0.85
@@
REVIDE Chen et al. [ChenHTYDK22] 20.10 0.85
TransWeather [Transweather] 17.33 0.82
WGWSNet [weather_data] 20.44 0.87
DSwinIR (Ours) 20.80 0.89
Average Chen et al. [ChenHTYDK22] 28.93 0.90
TransWeather [Transweather] 26.71 0.86
WGWSNet [weather_data] 29.61 0.90
DSwinIR (Ours) 30.51 0.92
TABLE VI: Quantitative comparison on the real-world WeatherBench dataset.
Type Method Venue Dehaze Derain Desnow Average
[…]
General DehazeFormer [DehazeFormer] TIP’23 24.12 0.745 0.345 36.05 0.954 0.181 28.88 0.849 0.178 29.68 0.849 0.235
DCMPNet [Zhang2024DCMPNet] CVPR’24 21.18 0.506 0.491 32.04 0.876 0.282 24.81 0.614 0.546 26.01 0.665 0.440
DRSformer [DRSformer] CVPR’23 19.95 0.694 0.404 33.98 0.943 0.209 28.00 0.836 0.197 27.31 0.824 0.270
NeRD-Rain [Chen2024nerd-rain] CVPR’24 21.52 0.718 0.386 35.74 0.950 0.182 28.87 0.851 0.182 28.71 0.840 0.250
SnowFormer [Chen2022Snowformer] arXiv’22 22.71 0.736 0.305 35.18 0.951 0.155 29.30 0.868 0.143 29.06 0.852 0.201
MPRNet [ZamirA2021Mprnet] CVPR’21 23.27 0.739 0.355 36.14 0.954 0.171 29.18 0.860 0.177 29.53 0.851 0.234
Restormer [Restormer] CVPR’22 19.30 0.687 0.412 34.49 0.945 0.197 27.95 0.836 0.197 27.25 0.823 0.269
All-in-One AirNet [AirNet] CVPR’22 20.94 0.705 0.383 33.59 0.942 0.224 22.06 0.780 0.291 25.53 0.809 0.299
TransWeather [Transweather] CVPR’22 19.79 0.680 0.397 29.34 0.903 0.294 24.96 0.796 0.231 24.70 0.793 0.307
WGWS-Net [weather_data] CVPR’23 13.79 0.603 0.535 37.08 0.961 0.117 20.81 0.780 0.248 23.89 0.781 0.300
PromptIR [PromptIR] NeurIPS’23 21.11 0.713 0.375 34.54 0.944 0.198 27.93 0.836 0.195 27.86 0.831 0.256
DiffUIR [Zheng2024DiffUIR] CVPR’24 22.74 0.744 0.355 35.93 0.955 0.172 29.50 0.870 0.162 29.39 0.856 0.230
MWFormer [Zhu2024MWFormer] TIP’24 24.42 0.746 0.284 35.15 0.951 0.153 29.98 0.872 0.133 29.85 0.856 0.190
Histoformer [Histoformer_SunRGWC24] ECCV’24 17.69 0.669 0.437 30.70 0.916 0.279 25.39 0.808 0.225 24.59 0.798 0.314
AdaIR [cui2025adair] ICLR’25 23.08 0.731 0.351 34.87 0.946 0.192 28.44 0.837 0.179 28.80 0.838 0.240
DSwinIR (Ours) 2025 25.15 0.766 0.329 37.17 0.961 0.173 30.12 0.873 0.170 30.81 0.867 0.224
Figure 7: Visual results on the real-world WeatherBench dataset.
@@
On a challenging benchmark composed of real-world rain, snow, and haze datasets [weather_data], DSwinIR again performs well. As shown in Table V, our model achieves an average PSNR of 30.51 dB, significantly outperforming the previous state-of-the-art WGWSNet [weather_data] by 0.9 dB.
Moreover, we conduct evaluations on the recent, large-scale real-world WeatherBench dataset [Guan2025WeatherBench], with results presented in Table VI. Our method, DSwinIR, achieves competitive overall performance. Notably, DSwinIR performs well in the Dehaze and Desnow tasks, outperforming the next-best method (MWFormer) by 0.73 dB in Dehaze PSNR and 0.14 dB in Desnow PSNR. While WGWS-Net shows strong performance in Derain, our model achieves a close second-best result. This performance across all categories results in a high overall average, indicating a good balance and robustness across diverse weather conditions. As illustrated in Figure 7, the qualitative results are consistent with these quantitative findings. DSwinIR effectively removes complex weather artifacts (e.g., dense fog, heavy snow) while preserving fine-grained details and avoiding color casts.
TABLE VII: Quantitative results on the Composite Degradation Setting.
@@
Restormer [Restormer] CVPR’22 26.29 28.35 33.10 33.43 24.80 25.28 24.99 26.80 26.15 23.90 23.82 26.99
NAFNet [chen2022simple] ECCV’22 24.50 25.34 29.24 29.54 21.91 22.75 22.79 23.67 23.86 21.03 20.82 24.13
AirNet [AirNet] CVPR’22 24.83 24.21 26.55 26.79 23.23 22.82 23.29 22.21 23.29 21.80 22.24 23.75
PromptIR [PromptIR] NeurIPS’23 26.32 26.10 31.56 31.53 24.49 25.05 24.51 24.54 23.70 23.74 23.33 25.90
TransWeather [Transweather] CVPR’22 23.39 23.95 26.69 25.74 22.24 22.62 21.80 23.10 22.34 21.55 21.01 23.13
WeatherDiff [weather_data] TPAMI’23 23.58 21.99 24.85 24.80 21.83 22.69 22.12 21.25 21.99 21.23 21.04 22.49
WGWSNet [weather_data] CVPR’23 24.39 27.90 33.15 34.43 24.27 25.06 24.60 27.23 27.65 23.90 23.97 26.96
InstructIR [InstructIR] ECCV’24 26.70 32.61 33.51 34.45 24.36 25.41 25.63 28.80 29.64 24.84 24.32 28.21
OneRestore [guo2024onerestore] ECCV’24 26.55 32.71 33.48 34.50 26.15 25.83 25.56 30.27 30.46 25.18 25.28 28.47
@@
IV-B3 Results on Composite Degradations
We evaluate DSwinIR on the challenging composite degradation benchmark (CDD-11) [guo2024onerestore], which features 11 combinations of low-light, haze, rain, and snow. As presented in Table VII, our DSwinIR achieves an average PSNR of 29.43 dB, which is 0.96 dB higher than the next-best method, OneRestore [guo2024onerestore]. The proposed DSwinIR can learn to adapt its receptive field to different degradation structures (e.g., thin rain streaks or broad haze), and our model achieves the advanced performance across all 11 individual and combined degradation settings, demonstrating the architecture’s robustness in handling complex, interacting artifacts. The qualitative results in Figure 8 are consistent with these quantitative findings, showing that DSwinIR effectively removes interacting artifacts while preserving image detail.
IV-B4 Results on Single‑Task Settings
To complement the all-in-one results, we also investigate the performance of DSwinIR on single-task benchmarks. These experiments are designed to assess the model’s effectiveness when its architecture is optimized for a single task, and to benchmark its performance against models developed specifically for these applications.
@@ IV-B4 Results on Single‑Task Settings
We first evaluate the model’s performance on the standard Rain100L benchmark. As shown in Table VIII, this is a highly competitive field. When trained solely for this task, DSwinIR demonstrates strong performance. The gain over other backbone architectures like PromptIR (+1.98 dB) suggests the architectural progress provided by our DSwin attention. Moreover, DSwinIR also achieves a 0.12 dB improvement over the recent AdaIR model.
TABLE VIII: Single-task deraining results on the Rain100L dataset.
Method DIDMDN [zhang2018DIDMDN] UMR [yasarla2019UMR] SIRR [wei2019sirr] MSPFN[jiang2020multi] LPNet[LPNet] AirNet[AirNet] Restormer[Restormer] PromptIR[PromptIR] AdaIR[cui2025adair] DSwinIR (Ours)
@@
The task of single-image dehazing presents a different challenge, often requiring a nuanced understanding of both global atmospheric light and local object details. As detailed in Table IX, DSwinIR achieves a PSNR of 32.89 dB. This result represents a 1.09 dB gain over the next-best method, AdaIR. We posit that the deformable mechanism’s ability to adapt its receptive field is well-suited to the non-homogeneous nature of haze, while the sliding mechanism’s capacity to preserve fine structural details ensures object clarity.
TABLE IX: Single-task dehazing results on the SOTS outdoor test set.
Method DehazeNet [DehazeNet] MSCNN [ren2016mscnn] AODNet [AODNet] EPDN [qu2019epdn] FDGAN [FDGAN] AirNet [AirNet] Restormer [Restormer] PromptIR [PromptIR] AdaIR [cui2025adair] DSwinIR (Ours)
@@
We also test DSwinIR on the real-world SPA-Data benchmark. The results, summarized in Table XII, show DSwinIR achieves a PSNR of 49.19 dB. This represents a 0.66 dB gain over the strong DRSformer model. It suggests that the DSwin attention’s ability to learn adaptive, data-driven receptive fields provides a practical advantage for tackling such complex, non-uniform degradations.
TABLE X: Quantitative comparison on the LOLv2 dataset [Yang2021LOLv2] for low-light enhancement.
Method LOLv2-Real LOLv2-Syn
PSNR SSIM $\uparrow$ PSNR $\uparrow$ SSIM
RetinexNet [Wei2018RetinexNet] 15.47 0.567 17.13 0.798
UFormer [Uformer] 18.82 0.771 19.66 0.871
FIDE [Xu2020FIDE] 16.85 0.678 15.20 0.612
DRBN [Yang2020DRBN] 20.29 0.831 23.22 0.927
KinD [Zhang2021KinD] 14.74 0.641 13.29 0.578
Restormer [Restormer] 19.94 0.827 21.41 0.830
MIRNet [Zamir2020MIRNet] 20.02 0.820 21.94 0.876
SNR-Net [Xu2022SNRNET] 21.48 0.849 24.14 0.928
FourLLIE [Wang2023FourLLIE] 21.60 0.847 24.17 0.917
Retinexformer [Retinexformer] 22.80 0.840 25.67 0.930
DSwinIR (Ours) 23.17 0.851 25.74 0.934
TABLE XI: Quantitative comparison for the mixed task (low-light enhancement and deblurring) on the LowBlur dataset [zhou2022lowblur].
Method KinD++ [Zhang2021KinD] DRBN [Yang2020DRBN] DeblurGAN-v2 [Kupyn2019Deblurgan] MIMO [Cho2021MIMO] NAFNet [chen2022simple] LEDNet [Zhou2022LEDNet] RetinexFormer [Retinexformer] Restormer [Restormer] DarkIR [Feijoo2025DarkIR] DSwinIR (Ours)
PSNR 21.26 21.78 22.30 22.41 25.36 25.74 26.02 26.72 27.30 27.33
SSIM 0.753 0.768 0.745 0.835 0.882 0.850 0.887 0.902 0.898 0.909
Low-Light Image Enhancement
We first evaluate DSwinIR on the widely-used low-light enhancement benchmark, LOLv2, which includes both real-world (LOLv2-Real) and synthetic (LOLv2-Syn) test sets. As presented in Table X, our model achieves state-of-the-art performance on both subsets. Furthermore, to assess the model’s capability in handling complex, compound degradations, we evaluate DSwinIR on the mixed task of joint low-light enhancement and deblurring. Following the experimental setup of DarkIR [Feijoo2025DarkIR] on the LowBlur dataset, our model again demonstrates superior performance, as detailed in Table XI. DSwinIR achieves the highest scores with 27.33 dB in PSNR and 0.909 in SSIM. These results underscore the strong generalization capability of our model for the dedicated low-light enhancement task.
TABLE XII: Single-task deraining results on the real-world SPA-Data dataset.
Method MSPFN [jiang2020multi] MPRNet [ZamirA2021Mprnet] DualGCN [li2021dual] SPDNet [huang2017spdnet] Uformer [Uformer] Restormer [Restormer] IDT [Xiao2023IDT] DRSformer [DRSformer] DSwinIR (Ours)
@@
TABLE XIII: Efficiency evaluation of different methods with $\times256$ input.
@@
AirNet [AirNet] Deformable Conv 9M 301G 0.57s 31.20
Uformer [Uformer] Window Attention 20M 41G 0.45s 30.66
Restormer [Restormer] Transposed Self-Attention 26M 155G 0.67s 30.75
PromptIR [PromptIR] Trans. SA + Prompts 36M 173G 0.71s 32.06
DRSformer [DRSformer] Sparse Window Attention 33M 243G 0.82s 31.26
@@ IV-C Ablation Study
To validate our design choices and rigorously quantify the contribution of each proposed component, we conduct a series of in-depth ablation studies. We use the challenging three-task all-in-one setting as our testbed and present the comprehensive results in Figure 9.
@@
To validate our design, we incrementally introduce our components starting from established baselines. The Uformer [Uformer] and DRSformer [DRSformer] models provide benchmarks at 30.66 dB and 31.26 dB, respectively. Our first experiment, isolating the sliding window mechanism, yields 31.79 dB, confirming the effectiveness of the token-centric paradigm. In parallel, isolating the deformable window mechanism achieves 32.11 dB, a result that already outperforms the prompt-based PromptIR baseline (32.06 dB). When these two components are combined, the resulting DSwin module (KK=7) achieves 32.53 dB. This performance gain is greater than the sum of the individual component improvements, demonstrating a clear synergistic effect.
Figure 9: Ablation studies demonstrating the effectiveness of our key components. We evaluate (1) different attention mechanisms, showing improvements from sliding window (31.79 dB) and Deformable Window (32.11 dB) over baselines; (2) DSwin configurations with various kernel sizes (KK=5, 7, 9) and multi-scale enhancement (32.69 dB); and (3) FFN module with MSG enhancement, achieving the best performance (32.72 dB). All experiments report the average PSNR of three distinct degradation tasks.
Figure 10: Visualization of receptive fields of multiple backbones by LAM[lam].
@@
The visualizations of the receptive field in Figure 10 offer profound qualitative insights into why DSwinIR is so effective. Its receptive field is demonstrably both global in reach and local in precision. This is the direct result of our design principles: the sliding mechanism ensures a connected, expansive base, while the deformable mechanism learns to dynamically concentrate attention on areas like object contours and textures.
Moreover, DSwinIR also exhibits computational efficiency, as detailed in Table XIII. It is crucial to note that DSwinIR achieves its superior performance without substantially increasing computational cost. Compared to its top-performing competitors, DSwinIR is significantly more economical. For instance, it requires 15% fewer FLOPs and is 22% faster than PromptIR, while delivering a substantially higher PSNR. The advantage is even more dramatic when compared to DRSformer, against which DSwinIR uses about 40% fewer FLOPs.
@@
Having validated the DSwin concept, we next investigate its optimal configuration. We first analyze the impact of a fixed kernel size, testing KK=5, KK=7, and KK=9. The results show a clear trend: performance improves from 32.02 dB (KK=5) to 32.53 dB (KK=7), and then begins to plateau at 32.61 dB (KK=9). This suggests a trade-off, where a larger receptive field is beneficial, but excessively large kernels may not provide proportionally greater returns for their computational cost. This observation motivates our multi-scale strategy. By employing multi-scale kernels within a single attention module (+Multi-Scale K), we reach a PSNR of 32.69 dB. This result is superior to even the best-performing single-kernel model (KK=9), providing clear evidence that empowering the model to simultaneously perceive features at multiple scales is more effective than relying on a single, large-but-fixed receptive field.
@@
Finally, we analyze the feed-forward network. Incorporating our proposed multi-scale gated FFN (+MSG) yields the model’s peak performance at 32.72 dB. While this incremental gain is modest, it confirms the benefit of maintaining a holistic multi-scale philosophy across the entire Transformer block.